OpenStudy (anonymous):

how much work did the movers do (horizontally) pushing a 160-kg crate 10.3 m across a rough floor without acceleration, if the effective coefficient of friction was 0.50?

OpenStudy (anonymous):

Work = F*d, the force due to friction is the force on the object perpindicular to the surface of friction (gravity acting on the box) times the coefficient of friction If the box is not accelerating, the force due to horizontal motion is 0, F=m*a where a=0, so the only force done on the box is friction, which is 160kg*9.8m/s^2*(.5) = 784N W=F*d => 784N*10.3m = 8075.2 N-m

OpenStudy (anonymous):

u still need help with this?